OpenStudy (anonymous):

Use the interative rule to find the 5th term of the sequence an = 3•4n-1 a5 =?????? PLEASE HELP

OpenStudy (anonymous):

♡★♡★♡★♡Welcome to OpenStudy! ♡★♡★♡★♡Please, don′t hesitate to ask questions!We′re here to help!⌣¨

OpenStudy (anonymous):

Yep ill help you

OpenStudy (landon34):

hey guys.

OpenStudy (anonymous):

Pleaseee help me

OpenStudy (landon34):

what is it?

OpenStudy (landon34):

oh.

OpenStudy (anonymous):

Please revoke the medal.

OpenStudy (anonymous):

someone please help me with my question

OpenStudy (anonymous):

I have been stuck on it for hours

OpenStudy (anonymous):

You need to insert 5, instead of n. \[a_n=3*4n-1\] \[a_5=3*4*5-1=?\]

OpenStudy (anonymous):

What does it equal?

Nnesha (nnesha):

the question is \[\rm a_n=3*4^{n-1}\] right ? where n= number of a term we need like to find 2nd term replace n with 2 to find 8th term replace n with 8

OpenStudy (anonymous):

wait then what do i do with the -1 ?

Nnesha (nnesha):

to find 5th term replace n with what ????

OpenStudy (anonymous):

So after i replace the n with a 5 what do i do ?

OpenStudy (anonymous):

Do i subtract 5-1?

Nnesha (nnesha):

correct!

OpenStudy (anonymous):

After that?

Nnesha (nnesha):

5-1 =4 so now simplify \[a_5=3*4^4\]

OpenStudy (anonymous):

a5 = 48??

OpenStudy (anonymous):

so i multiply 3 • 16??

Nnesha (nnesha):

well what's the question is it 4 to the (n-1) power ??? if yes then no 48 isn't correct

OpenStudy (anonymous):

then do i multiply 4•4•4•4 ??????

Nnesha (nnesha):

correct the exponent represent how many times we should multiply the base by itself

OpenStudy (anonymous):

Then when i get that answer i multiply it by 3?

OpenStudy (anonymous):

I got 3•256 = 768 SO A5=768?????

OpenStudy (anonymous):

Hello?

OpenStudy (anonymous):

I just want to make sure

Nnesha (nnesha):

yes that's correct . and sorry for late reply